MOA-II Obsevation Log Monthly Summary 2007/08

Return to index

Previous month Next month

TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB9412-102468334450260950
LMC1610-169283481140460
SMC415-43218349960
ALERT-000
GRB-000
TOTAL---9344965411370 (114h16m10s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
177161144144646131145217665127171137
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
1441221181281192411521771772070


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
182019181820828072292226
Field13Field14Field15Field16
292800


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
17170000000000


SPECIAL TARGET NAME

CategoryTargetNrun


CCD TEMPERATURE Previous month Next month




VACUUM Previous month Next month




TELESCOPE TEMPERATURE Previous month Next month




COMMENTS
2007/08/1: DARK D2521(60sec) and D2522(300sec) are OK. / night time = 44220(sec)

2007/08/2: DARK D2524(60sec) and D2525(300sec) are OK. / night time = 44100(sec)

2007/08/3: DARK D2526(60sec) and D2527(60sec) are OK.. / night time = 43980(sec)

2007/08/4: DARK D2528(60sec) is OK. But no observation. / night time = 43860(sec)

2007/08/5: DARK D2532(60sec) is OK. Power cut between 20:40-22:16. / night time = 43740(sec)

2007/08/6: DARK D2539(60sec) is OK. / night time = 43620(sec)

2007/08/7: DARK D2541(60sec) and D2542 are OK. / night time = 43500(sec)

2007/08/8: Dark D2544(60sec), D2546(60sec) are OK. Seeing is too bad (more than 4 arcsec). / night time = 43380(sec)

2007/08/9: Dark 2549(60sec) is OK. CCD temparture becames nearly -60 degree, then we start the pumping at am 2:05 / night time = 43260(sec)

2007/08/10: Very strong wind, so that we test the auto shout down at the Power Failure, and test for filter senser. / night time = 43140(sec)

2007/08/11: Dark2556(300sec) is OK. Due to the Frost , start is delay. / night time = 42960(sec)

2007/08/12: Dark D2562(60sec),D2563(60sec) are OK. / night time = 42840(sec)

2007/08/13: Dark D2567(60sec), D2566 (300sec) are OK. / night time = 42720(sec)

2007/08/14: Some of the images are trailing in left side and rightside. This cause due to shutter close time is longer than expected. Change the setting the moa_cam3 by KOKI, it works well. Dark D2570(60sec) for GB B9676-B9700. Dark D2573(60sec) for GB B9701-B9724. Dark D2572(300sec) is OK. / night time = 42600(sec)

2007/08/15: D2576(300sec) and D2577(60sec) are OK. / night time = 42420(sec)

2007/08/16: Dark D2579(60sec) D2580(60sec) are OK. Peak of MOA373(gb13) is followup. After the midnight. vacuum become worse, so we start the pumping at AM 2:00. / night time = 42300(sec)

2007/08/17: Dark D2587(60sec),D2588(60sec) and D2589(300sec) are OK. / night time = 42120(sec)

2007/08/18: Dark D2591(60sec) for B9806-9808. Dark D2600(60sec) and D2602(60sec) are for B9809-B9814. Dark D2601(300sec) is OK. / night time = 42000(sec)

2007/08/19: Dark D2609(60sec),D2610(60sec) and D2608(300sec) are OK. High magnification microlensing MOA-378 is in gb8-chip7. / night time = 41880(sec)

2007/08/20: Dark D2615-a,b (300sec) D2616-a(300sec), and D2613(60sec) are OK. / night time = 41700(sec)

2007/08/21: Anomaly on MOA-365 (gb18). Dark D2620(300sec) and D2621(60sec) are OK. / night time = 41580(sec)

2007/08/22: Dark D2625(300sec) and D2626(60sec) are OK. / night time = 41400(sec)

2007/08/23: Dark D2630(300sec) and D2631(60sec) are OK. Only fields gb19,20,21,22 are observed due to Moon. / night time = 41280(sec)

2007/08/24: Dark D2634(300sec), D2635(40sec) and D2636(60sec) are OK. Moon is close to the Galactic Bulge. B10042-B10046 are 60 sec exp. B10047-B10096 are 40 sec exp. MOA365(gb18) is anomaly. / night time = 41100(sec)

2007/08/25: Dark D2639(40sec) and D2640(60sec) are OK. / night time = 40920(sec)

2007/08/26: Dark D2646(60sec) and D2645(300sec) are OK. / night time = 40800(sec)

2007/08/27: Dark D2650(60sec) andD2649(300sec) are OK. / night time = 40620(sec)

2007/08/28: Dark D2653(180sec) and D2654(60sec) are OK. Lunar Eclipes from pm 21h to am 0h. / night time = 40440(sec)

2007/08/29: D2658(15sec) for FLAT is OK / night time = 40320(sec)

2007/08/30: D2659(60sec) and D2660(300sec) are OK / night time = 40140(sec)

2007/08/31: No observation / night time = 39960(sec)

TOTAL NIGHT TIME: 1308900(sec)

OPERATING RATE: 46.6 %
GB: 33.5 %
LMC: 12.2 %
SMC: 0.9 %
ALERT: 0.0 %
GRB: 0.0 %

Return to index

Previous month Next month